Sarah and Rachel are sisters. They leave a park at the same time on their bycicle and ride home along the same bycicle path. Sarah is in a hurry, so she cycles at 15km/h. Rachel has time to spare, so she cycles at 11km/h. Sarah gets home at 12 min before Rachel. How long did Sarah take to ride home from the park?
Answer by stanbon(75887)

Sarah and Rachel are sisters. They leave a park at the same time on their bycicle and ride home along the same bycicle path. Sarah is in a hurry, so she cycles at 15km/h. Rachel has time to spare, so she cycles at 11km/h. Sarah gets home at 12 min before Rachel. How long did Sarah take to ride home from the park?
-------
Sarah DATA:
rate = 15 km/h ; distance = x km ; time = x/15 hrs
----
Rachel DATA:
rate = 11 km/h ; distance = x km ; time = x/11 hrs
-----
Equation:
rachel time - sarah time = 12 min = 1/5 hr
-----
x/11 - x/15 = 1/5
-----
75x - 55x = 11*15
20x = 156
x = 7.8 km
--------
Sarah time = x/15 = 7.8/15 = 0.52 hr = 31.2 min = 31min 12sec
=================================================================
Cheers,
Stan H.
==============